hi
i am studying for my final and i would like to know something. this is the task:
tamika invests $3000 in a 5-yr cd that pays 5.35% interest compounded monthly. how much will she have in the acc. at the end of the 5 yr period?
I used the formula A=P(1+r/n)nt
so I got A = 3000(1+0.535/12)60 (that is an exponent)
now, i first divided 0.535 by 12 and then +1 so I get 1.044583333... and i took that by 60 exponential. I get 13.69574101
on the answer sheet it is written like this:
A=3000(1+0.535/12)60
3000(1.305916985) = 3917.75
how do i get the 1.3059...etc? how do i calculate the braces with the 60 exponential part? what i did was calculating the braces first and then take it exponential 60.

Answer by stanbon(57250)

tamika invests $3000 in a 5-yr cd that pays 5.35% interest compounded monthly. how much will she have in the acc. at the end of the 5 yr period?
I used the formula A=P(1+r/n)nt
so I got A = 3000(1+0.535/12)60 (that is an exponent)
now, i first divided 0.535 by 12 and then +1 so I get 1.044583333... and i took that by 60 exponential. I get 13.69574101
on the answer sheet it is written like this:
A=3000(1+0.535/12)60
3000(1.305916985) = 3917.75
how do i get the 1.3059...etc? how do i calculate the braces with the 60 exponential part? what i did was calculating the braces first and then take it exponential 60.
----------------------
Your mistake is changing 5.35% to 0.535.
You should have 5.35% = 0.0535
----------------------
After that your procedure is correct.
----------------------
Cheers,
Stan H.